# Global convergence of a modified BFGS-type method based on function information for nonconvex multiobjective optimization problems

### v1

@@ 2 Preliminaries
$x^{\top}B_{k+1}x\geq\dfrac{(x^{\top}\gamma_{k})^{2}}{\gamma_{k}^{\top}s_{k}}\geq 0.$ (4.12)
Inequality 4.12 holds as an equality when $x^{\top}\gamma_{k}=0$ and $x=cs^{k}$, where $c$ is a constant. It follows that $cs_{k}^{\top}\gamma_{k}=0$. By (4.10), we can obtain that $c=0$, which contradicts the fact that $x$ is a nonzero vector.
@@
For any $z\in R^{n}$, the level set $\left\{x:f(x)\leq f(z)\right\}$ is a bounded set.
@@
The proof is similar to that in [21, Proposition 2], we omit it here. ∎
@@
$(i)$ The point $x\in{{R}^{n}}$ is not Pareto critical;
@@
$(c)$ If there exists $K\subset N$ such that $x_{k}\stackrel{{\scriptstyle K}}{{\longrightarrow}}x^{*}$, $d_{k}\stackrel{{\scriptstyle K}}{{\longrightarrow}}0$, then $x^{*}$ is a Pareto critical point.
@@
These assertions can be derived by using the same proof demonstrated in the [32, Proposition 4.1]. ∎
@@ 4.1 Global convergence
(ii)For any $x\in R^{n}$, there exists constant $a$ and $b$ such that
@@
In the subsection, the following properties imply that under the Wolfe line search, the objective function will be significantly reduced along the direction generated by Algorithm 1.
@@
Let $B_{k}$ be generated by the BFGS update formula (4.6) in Algorithm 1. If there exists constants $c_{1}$ and $c_{2}$ such that $\frac{\gamma_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}\geq c_{1}>0$, $\frac{\|\gamma_{k}\|^{2}}{\gamma_{k}^{\top}s_{k}}\leq c_{2}$. Then, for any $p\in(0,1)$, there exists a constant $\delta>0$ such that, for any $k\in N$,
$\cos\beta_{j}=\frac{s_{j}^{\top}B_{j}s_{j}}{\|s_{j}\|\|B_{j}s_{j}\|}\geq\delta,$ (4.13)
holds for at least $\lceil pk\rceil$ values of $j\in[1,k].$
@@
The proof is similar to that in [22, Theorem 2.1], we omit it here. ∎
@@
$\frac{\gamma_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}\geq c_{1}\|d_{k}\|_{B_{k}}^{2},$ (4.14)
$\frac{\|\gamma_{k}\|^{2}}{\gamma_{k}^{\top}s_{k}}\leq\frac{c_{2}}{\|d_{k}\|_{B_{k}}^{2}}.$ (4.15)
@@
First, inequality (4.14) can be deduced from Wolfe line search, (3.10) and (4.9), i.e.,
$f_{i}(x_{k})-f_{i}(x_{k+1})\geq-\sigma_{1}\alpha_{k}\mathcal{D}(x_{k},d_{k})\geq\sigma_{1}\alpha_{k}\|d_{k}\|_{B_{k}}^{2},$
$\gamma_{k}^{\top}s_{k}\geq\sum_{i=1}^{m}\lambda_{i}^{k}\left(f_{i}(x^{k})-f_{i}(x^{k+1})\right)\|s_{k}\|^{2}.$
We denote $c_{1}=\sigma_{1}\alpha_{1}>0$, it follows by Proposition 4.1 that (4.14) holds. Next, we consider that inequality (4.15). By the definition of the 4.5,
$\|\gamma_{k}\|$ $\leq\|y_{k}\|+m_{k}\|s_{k}\|$
$\leq\left(\frac{\|y_{k}\|}{\|s_{k}\|}+m_{k}\right)\|s_{k}\|$
$\leq\left[\frac{\|y_{k}\|}{\|s_{k}\|}+\max\{-\eta_{k},0\}+\sum_{i=1}^{m}\lambda_{i}^{k}\left(f_{i}(x^{k})-f_{i}(x^{k+1})\right)\right]\|s_{k}\|.$ (4.16)
@@
$\frac{\|y_{k}\|}{\|s_{k}\|}\leq\sum_{i\in[m]}\lambda_{i}^{k}\frac{\|\nabla f_{i}(x_{k+1})-\nabla f_{i}(x_{k})\|}{\|x_{k+1}-x_{k}\|}\leq\max_{i\in[m]}L_{i}.$
@@
$\frac{|\langle y_{k},s_{k}\rangle|}{\|s_{k}\|^{2}}\leq\frac{\|y_{k}\|}{\|s_{k}\|}.$
@@
$\frac{\|y_{k}\|}{\|s_{k}\|}+\max\{-\eta_{k},0\}\leq\frac{2\|y_{k}\|}{\|s_{k}\|}\leq 2\max_{i\in[m]}L_{i}.$ (4.17)
Assumption 4.1 and Assumption 4.2 (i) imply the existence of $f_{1}\in R$ such that $f_{i}(x_{k})\geq f_{1}$ for all $k\geq 0$ and $i\in[m]$. It follows by (4.1) and (4.17) that
$\|\gamma_{k}\|\leq(2\max_{i\in[m]}L_{i}+\max_{i\in[m]}(f_{i}(x_{0})-f_{1}))\|s_{k}\|,$
@@
$\frac{\|\gamma_{k}\|^{2}}{\gamma_{k}^{\top}s_{k}}\leq\frac{c_{2}}{\|d_{k}\|_{B_{k}}^{2}},$
where $c_{2}=\frac{(2\max_{i\in[m]}L_{i}+\max_{i\in[m]}(f_{i}(x_{0})-f_{1}))^{2}}{c_{1}}.$
@@
$\sum_{k\geq 0}\frac{\mathcal{D}(x_{k},d_{k})}{\|d_{k}\|^{2}}<\infty.$ (4.19)
@@
The proof is similar to that in [14, Proposition 3.3], we omit it here. ∎
@@
$\|d_{SD}^{k}\|\geq\varepsilon,\quad\forall k\in N.$ (4.20)
By the definition of $d_{k}$ and Assumption 4.2 (iii), we get
$-\frac{1}{2}\|d_{k}\|_{B_{k}}^{2}$ $=\min\limits_{d\in R^{n}}\max\limits_{i\in[m]}{\nabla f_{i}(x_{k})^{\top}d+\frac{1}{2}}\|d\|_{B_{k}}^{2}$
$\leq\min\limits_{d\in R^{n}}\max\limits_{i\in[m]}{\nabla f_{i}(x_{k})^{\top}d+\frac{b}{2}}\|d\|^{2}$
$\leq\frac{1}{b}\min\limits_{d\in R^{n}}\max\limits_{i\in[m]}{\nabla f_{i}(x_{k})^{\top}(bd)+\frac{1}{2}}\|bd\|^{2}$
@@
$\frac{\gamma_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}\geq C_{1}>0,$
$\frac{\|\gamma_{k}\|^{2}}{\gamma_{k}^{\top}s_{k}}\leq C_{2}.$
Then, by Lemma 4.3 and Lemma 4.2, there exist a constant $\delta>0$ and a infinite subsequence of indices $\mathcal{K}\subset N$ such that
$-\frac{\mathcal{D}(x_{k},d_{k})}{\|d_{k}\|}\geq\delta\|d_{SD}^{k}\|\geq\delta\varepsilon,\quad\forall k\in\mathcal{K},$
@@
$\sum\limits_{k\in N}\frac{\mathcal{D}(x_{k},d_{k})^{2}}{\|d_{k}\|^{2}}\geq\sum\limits_{k\in\mathcal{K}}\frac{\mathcal{D}(x_{k},d_{k})^{2}}{\|d_{k}\|^{2}}\geq\sum\limits_{k\in\mathcal{K}}\delta^{2}\varepsilon^{2}=\infty,$
@@
Assumption 4.1 implies there exists a infinite subsequence of indices $K_{1}\subset{N}$ such that $x_{k}\stackrel{{\scriptstyle K_{1}}}{{\longrightarrow}}x^{*}$ and $\|d_{SD}^{k}\|\stackrel{{\scriptstyle K_{1}}}{{\longrightarrow}}0.$ By the Lemma 3.1, $d_{SD}(x)$ is continuous, we can get
@@
Furthermore, $x^{*}$ is Pareto critical. ∎
@@ 4.2 Superlinear convergence analysis
(ii) For $i\in[m]$, $\{\gamma_{k}\}$ satisfies $\sum\limits_{k>0}\gamma_{k}<\infty.$
(iii) There exists constants $U,L>0$ such that $UI_{n}\preceq\nabla^{2}f_{i}(x)\preceq LI_{n}.$
(iv) $\nabla^{2}f_{i}(x)$ is Lipschitz continuous with constant $M$, $i\in[m].$
(v)For any $x\in R^{n}$, there exists constant $a$ and $b$ such that
@@
The Assumption 4.3 $(iii)$ implies that $f_{i}(x)$ is strongly convex with modulus $U$ and $\nabla^{2}f_{i}(x)$ is Lipschitz continuous with constant $L$, for $i\in[m]$. In the previous section, we obtain the weak convergence of the sequence $\{x^{k}\}$ generated by Algorithm 1. Under strict convexity, the following theorem demonstrates that the sequence $\{x^{k}\}$ generated by Algorithm 1 strongly converges to the Pareto optimal point $x^{*}.$
Theorem 4.3.
Let $\{x_{k}\}$ be the sequence generated by Algorithm 1. Suppose that $f(x)$ is strictly convex, Assumption 4.1 and Assumption 4.2 hold. Then, $\{x_{k}\}$ converges to Pareto optimal point $x^{*}$ of (MOP).
Proof.
The proof is similar to that in [20, Theorem 3.6], we omit it here. ∎
@@
Suppose that Assumption 4.3 holds and $\{x_{k}\}$ is the sequence generated by Algorithm 1. Then, for any $k\geq 0$,
(i) $\|x_{k}-x^{*}\|\leq\frac{2}{U}\|d_{SD}^{k}\|,$
(ii) $\|s_{k}\|\geq\frac{1-\sigma_{2}}{L}\cos\beta_{k}\|d_{SD}^{k}\|.$
@@
The two assertions can be obtained by using the same arguments as in the proof of [19, Lemma 4.4], we omit them here. ∎
@@
Let $\left\{{{x}_{k}}\right\}$ be the sequence generated by Algorithm 1. Suppose that Assumption 4.3 holds. Then, there exists constant $r$ such that
$\frac{\gamma_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}\geq U,$ (4.21)
$\frac{\|\gamma_{k}\|^{2}}{\gamma_{k}^{\top}s_{k}}\leq\frac{r^{2}}{U}.$ (4.22)
@@
$\frac{\gamma_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}=\frac{(y_{k}+m_{k}s_{k})^{\top}s_{k}}{\|s_{k}\|^{2}}=\frac{y_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}+m_{k}\geq U.$ (4.23)
Since each objective function is U-strongly convex, for $i\in[m]$ and any $x\in{{R}^{n}}$, ${{f}_{i}}(x)$ has a lower bound $f_{i}^{\min}.$
@@
$\|\gamma_{k}\|\leq r\|s_{k}\|,$
where $r=2\max\limits_{i\in[m]}L_{i}+\max\limits_{i\in[m]}(f_{i}(x_{0})-f_{i}^{min}).$
@@
Suppose that Assumption 4.3 holds and $\{x_{k}\}$ is the sequence generated by Algorithm 1. Let $x^{*}$ be defined as Theorem 4.3. Then,
@@
The proof of the theorem is similar to the corresponding discussion in [21, Proposition 4.1], so we omit the proof here. ∎
@@
Using the proof of Theorem 4.4, it can directly deduce that a constant $r\in(0,1)$ exists such that
$0<\sum\limits_{i\in[m]}\lambda_{i}^{k}(f_{i}(x_{k})-f_{i}(x_{k+1}))<r^{k},\ \forall k\geq 0,$
where $\lambda_{k}={{\left(\lambda_{1}^{k},\lambda_{2}^{k},\dots,\lambda_{m}^{k}\right)}^{T}}$is the solution of (3.9). In the strongly convex case, we can observe that $\eta_{k}>0$ and
@@
$m_{k}$ $=\max\{-\eta_{k},0\}+\sum_{i=1}^{m}\lambda_{i}^{k}(f_{i}(x_{k})-f_{i}(x_{k+1}))$
$=\sum_{i=1}^{m}\lambda_{i}^{k}(f_{i}(x_{k})-f_{i}(x_{k+1}))\to 0,\ k\to\infty.$
in (4.5). In this case, $B_{k+1}s_{k}\approx y_{k}$. Then (4.8) can be seen as an approximation of the secant equation in [15]. Remarkably, for strongly convex functions, both MFONMO and MQNMO in [15] may exhibit comparable convergence rates. This observation is supported by the numerical experiments presented in the following section.
@@
$f_{\lambda}(x)=\sum\limits_{i\in[m]}\lambda_{i}^{k}f_{i}(x),$
@@
where $\lambda_{k}={{\left(\lambda_{1}^{k},\lambda_{2}^{k},\dots,\lambda_{m}^{k}\right)}^{T}}$is the solution of (3.9). Let us define
$\bar{s}_{k}=\nabla^{2}f_{\lambda}(x^{*})^{\frac{1}{2}}s_{k},\quad\bar{y}_{k}=\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}y_{k},\quad\bar{\gamma}_{k}=\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}\gamma_{k},$
$\bar{B}_{k}=\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}B_{k}\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}},$ (4.25)
@@
Suppose that Assumption 4.3 holds. Let $\{x_{k}\}$ be the sequence generated by Algorithm 1. Then
$\lim\limits_{k\to\infty}\frac{\|(B_{k}-\nabla^{2}f_{\lambda}(x^{*}))s_{k}\|}{\|s_{k}\|}=0.$
@@
$(\bar{G}_{\lambda}^{k}-\nabla^{2}f_{\lambda}(x^{*}))s_{k}=y^{k}-\nabla^{2}f_{\lambda}(x^{*})s_{k},\quad k\geq 0.$ (4.27)
It follows by Assumption 4.3 (iv), (4.27) and the definition of $\bar{y}_{k}$, $\bar{s}_{k}$ that
$\|\bar{y}_{k}-\bar{s}_{k}\|$ $=\|\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}(\bar{G}_{\lambda}^{k}-\nabla^{2}f_{\lambda}(x^{*}))\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}\bar{s}_{k}\|$
$\leq\|\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}\|^{2}\|\bar{s}_{k}\|\|\bar{G}_{\lambda}^{k}-\nabla^{2}f_{\lambda}(x^{*})\|$
@@
$\leq\bar{c}\|\bar{s}_{k}\|\varepsilon_{k},$ (4.28)
where $\bar{c}=M\|\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}\|^{2}$, $\varepsilon_{k}=\max\{\|x_{k+1}-x^{*}\|,\|x_{k}-x^{*}\|\}.$ Since
$|\|\bar{y}_{k}\|-\|\bar{s}_{k}\||\leq\|\bar{y}_{k}-\bar{s}_{k}\|,$

### v2

@@ 2 Preliminaries
$x^{\top}B_{k+1}x\geqslant\dfrac{(x^{\top}\gamma_{k})^{2}}{\gamma_{k}^{\top}s_{k}}\geqslant 0.$ (4.12)
Since $x^{\top}B_{k+1}x=0$ if and only if $x^{\top}\gamma_{k}=0$ and $x=cs^{k}$, where $c$ is a constant. It follows that $cs_{k}^{\top}\gamma_{k}=0$. By (4.10), we can obtain that $c=0$ and $x=0.$ Then for all nonzero vector $x\in\mathbb{R}^{n},$
$x^{\top}B_{k+1}x>0,$
we conclude the proof. ∎
@@
For any $z\in\mathbb{R}^{n}$, the level set $\left\{x:f(x)\leqslant f(z)\right\}$ is a bounded set.
@@
The proof is similar to that in [14, Proposition 2], we omit it here. ∎
@@
$(i)$ The point $x\in{{\mathbb{R}}^{n}}$ is not Pareto critical;
@@
These assertions can be derived by using the same proof demonstrated in the [25, Lemma 2.2]. ∎
@@ 4.1 Global convergence
(ii)For any $x\in\mathbb{R}^{n}$, there exists constant $a$ and $b$ such that
@@
In the subsection, the following properties imply that under the Wolfe line search, the objective function will be significantly decreased along the direction generated by Algorithm 1.
@@
Let $B_{k}$ be generated by the BFGS update formula (4.6) in Algorithm 1. If there exists constants $c_{1}$ and $c_{2}$ such that $\frac{\gamma_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}\geqslant c_{1}>0$, $\frac{\|\gamma_{k}\|^{2}}{\gamma_{k}^{\top}s_{k}}\leqslant c_{2}$. Then, for any $p\in(0,1)$, there exists a constant $\delta>0$ such that, for any $k\in\mathbb{N}$,
$\cos\beta_{j}=\frac{s_{j}^{\top}B_{j}s_{j}}{\|s_{j}\|\|B_{j}s_{j}\|}\geqslant\delta,$ (4.13)
holds for at least $\lceil p(k+1)\rceil$ values of $j\in[0,k].$
@@
The proof is similar to that in [4, Theorem 2.1], we omit it here. ∎
@@
$\frac{\gamma_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}\geqslant c_{1}\|d_{k}\|_{B_{k}}^{2},$ (4.14)
$\frac{\|\gamma_{k}\|^{2}}{\gamma_{k}^{\top}s_{k}}\leqslant\frac{c_{2}}{\|d_{k}\|_{B_{k}}^{2}}.$ (4.15)
@@
From Wolfe line search, (3.10) and (4.9), we can get
$f_{i}(x_{k})-f_{i}(x_{k+1})\geqslant-\sigma_{1}\alpha_{k}\mathcal{D}(x_{k},d_{k})\geqslant\sigma_{1}\alpha_{k}\|d_{k}\|_{B_{k}}^{2},$
$\gamma_{k}^{\top}s_{k}\geqslant\sum_{i=1}^{m}\lambda_{i}^{k}\left(f_{i}(x_{k})-f_{i}(x_{k+1})\right)\|s_{k}\|^{2}.$
Let $c_{1}=\sigma_{1}\alpha_{1}>0$, it follows by Proposition 4.1 that
$\frac{\gamma_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}\geqslant c_{1}\|d_{k}\|_{B_{k}}^{2}.$
Next, we consider that inequality (4.15). By the definition of the 4.5,
$\|\gamma_{k}\|$ $\leqslant\|y_{k}\|+m_{k}\|s_{k}\|$
$\leqslant\left(\frac{\|y_{k}\|}{\|s_{k}\|}+m_{k}\right)\|s_{k}\|$
$\leqslant\left[\frac{\|y_{k}\|}{\|s_{k}\|}+\max\{-\eta_{k},0\}+\sum_{i=1}^{m}\lambda_{i}^{k}\left(f_{i}(x_{k})-f_{i}(x_{k+1})\right)\right]\|s_{k}\|.$ (4.16)
@@
$\frac{\|y_{k}\|}{\|s_{k}\|}\leqslant\sum_{i\in[m]}\lambda_{i}^{k}\frac{\|\nabla f_{i}(x_{k+1})-\nabla f_{i}(x_{k})\|}{\|x_{k+1}-x_{k}\|}\leqslant\max_{i\in[m]}L_{i}.$
@@
$\frac{|\langle y_{k},s_{k}\rangle|}{\|s_{k}\|^{2}}\leqslant\frac{\|y_{k}\|}{\|s_{k}\|}.$
@@
$\frac{\|y_{k}\|}{\|s_{k}\|}+\max\{-\eta_{k},0\}\leqslant\frac{2\|y_{k}\|}{\|s_{k}\|}\leqslant 2\max_{i\in[m]}L_{i}.$ (4.17)
Assumption 4.1 and Assumption 4.2 (i) imply the existence of $f_{1}\in\mathbb{R}$ such that $f_{i}(x_{k})\geqslant f_{1}$ for all $k\geqslant 0$ and $i\in[m]$. It follows by (4.1) and (4.17) that
$\|\gamma_{k}\|\leqslant(2\max_{i\in[m]}L_{i}+\max_{i\in[m]}(f_{i}(x_{0})-f_{1}))\|s_{k}\|,$
@@
$\frac{\|\gamma_{k}\|^{2}}{\gamma_{k}^{\top}s_{k}}\leqslant\frac{c_{2}}{\|d_{k}\|_{B_{k}}^{2}},$
where $c_{2}=\frac{(2\max\limits_{i\in[m]}L_{i}+\max\limits_{i\in[m]}(f_{i}(x_{0})-f_{1}))^{2}}{c_{1}}.$
@@
$\sum_{k\geqslant 0}\frac{\mathcal{D}(x_{k},d_{k})^{2}}{\|d_{k}\|^{2}}<\infty.$ (4.19)
@@
The proof is similar to that in [15, Proposition 3.3], we omit it here. ∎
@@
$\|d_{SD}^{k}\|\geqslant\varepsilon,\quad\forall k\geqslant 0.$ (4.20)
By the definition of $d_{k}$ and Assumption 4.2 (ii), we get
$-\frac{1}{2}\|d_{k}\|^{2}$ $=\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{i\in[m]}{\nabla f_{i}(x_{k})^{\top}d+\frac{1}{2}}\|d\|_{B_{k}}^{2}$
$\leqslant\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{i\in[m]}{\nabla f_{i}(x_{k})^{\top}d+\frac{b}{2}}\|d\|^{2}$
$\leqslant\frac{1}{b}\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{i\in[m]}{\nabla f_{i}(x_{k})^{\top}(bd)+\frac{1}{2}}\|bd\|^{2}$
@@
$\frac{\gamma_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}\geqslant C_{1}>0,$
$\frac{\|\gamma_{k}\|^{2}}{\gamma_{k}^{\top}s_{k}}\leqslant C_{2}.$
Then, by Lemma 4.3 and Lemma 4.2, there exist a constant $\delta>0$ and a infinite subsequence of indices $\mathcal{K}\subset\mathbb{N}$ such that
$-\frac{\mathcal{D}(x_{k},d_{k})}{\|d_{k}\|}\geqslant\delta\|d_{SD}^{k}\|\geqslant\delta\varepsilon,\quad\forall k\in\mathcal{K},$
@@
$\sum\limits_{k\in\mathbb{N}}\frac{\mathcal{D}(x_{k},d_{k})^{2}}{\|d_{k}\|^{2}}\geqslant\sum\limits_{k\in\mathcal{K}}\frac{\mathcal{D}(x_{k},d_{k})^{2}}{\|d_{k}\|^{2}}\geqslant\sum\limits_{k\in\mathcal{K}}\delta^{2}\varepsilon^{2}=\infty,$
@@
Assumption 4.1 implies there exists a infinite subsequence of indices $K_{1}\subset{\mathbb{N}}$ such that $x_{k}\stackrel{{\scriptstyle K_{1}}}{{\longrightarrow}}x^{*}$ and $\|d_{SD}^{k}\|\stackrel{{\scriptstyle K_{1}}}{{\longrightarrow}}0.$ By the Lemma 3.1, $d_{SD}(x)$ is continuous, we can get
@@
Furthermore, by Lemma 4.1, $x^{*}$ is Pareto critical. ∎
We have obtained the weak convergence of the sequence $\{x^{k}\}$ generated by Algorithm 1. Under strict convexity, the following theorem demonstrates that the sequence $\{x^{k}\}$ generated by Algorithm 1 strongly converges to the Pareto optimal point $x^{*}.$
Theorem 4.3.
Let $\{x_{k}\}$ be the sequence generated by Algorithm 1. Suppose that $f(x)$ is strictly convex, Assumption 4.1 and Assumption 4.2 hold. Then, $\{x_{k}\}$ converges to a Pareto optimal point $x^{*}$ of (MOP).
Proof.
The proof is similar to that in [25, Theorem 3.6], we omit it here. ∎
@@ 4.2 Superlinear convergence analysis
(ii) There exists constants $U,L>0$ such that $UI_{n}\preceq\nabla^{2}f_{i}(x)\preceq LI_{n},$ for $i\in[m].$
(iii) $\nabla^{2}f_{i}(x)$ is Lipschitz continuous with constant $M$, $i\in[m].$
(iv) For any $x\in\mathbb{R}^{n}.$ there exists constant $a$ and $b$ such that
@@
The Assumption 4.3 $(ii)$ implies that $f_{i}(x)$ is strongly convex with modulus $U$ and $\nabla^{2}f_{i}(x)$ is Lipschitz continuous with constant $L$, for $i\in[m]$.
@@
Suppose that Assumption 4.1 and Assumption 4.3 hold. Let $\{x_{k}\}$ be the sequence generated by Algorithm 1. Then, for any $k\geqslant 0$,
(i) $\|x_{k}-x^{*}\|\leqslant\frac{2}{U}\|d_{SD}^{k}\|,$
(ii) $\|s_{k}\|\geqslant\frac{1-\sigma_{2}}{L}\cos\beta_{k}\|d_{SD}^{k}\|.$
@@
The two assertions can be obtained by using the same arguments as in the proof of [26, Lemma 4.4], we omit them here. ∎
@@
Let $\left\{{{x}_{k}}\right\}$ be the sequence generated by Algorithm 1. Suppose that Assumption 4.1 and Assumption 4.3 hold. Then, there exists a constant $r$ such that
$\frac{\gamma_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}\geqslant U,$ (4.21)
$\frac{\|\gamma_{k}\|^{2}}{\gamma_{k}^{\top}s_{k}}\leqslant\frac{r^{2}}{U}.$ (4.22)
@@
$\frac{\gamma_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}=\frac{(y_{k}+m_{k}s_{k})^{\top}s_{k}}{\|s_{k}\|^{2}}=\frac{y_{k}^{\top}s_{k}}{\|s_{k}\|^{2}}+m_{k}\geqslant U.$ (4.23)
Since each objective function is U-strongly convex, for $i\in[m]$ and any $x\in{{\mathbb{R}}^{n}}$, ${{f}_{i}}(x)$ has a lower bound $f_{i}^{\min}.$
@@
$\|\gamma_{k}\|\leqslant r\|s_{k}\|,$
where $r=2\max\limits_{i\in[m]}L_{i}+\max\limits_{i\in[m]}(f_{i}(x_{0})-f_{i}^{min})$, the proof is complete.
@@
Suppose that Assumption 4.1 and Assumption 4.3 hold. Let $\{x_{k}\}$ be the sequence generated by Algorithm 1 and $x^{*}$ is defined as Theorem 4.3. Then,
@@
The proof of the theorem is similar to the corresponding discussion in [14, Proposition 4.1], so we omit the proof here.
@@
Using the convexity of $f(x)$, Proposition 4.1 and Assumption 4.3(iv), we can get
$0\leqslant\sum\limits_{i\in[m]}\lambda_{i}^{k}(f_{i}(x_{k})-f_{i}(x_{k+1}))$ $\leqslant(\sum\limits_{i\in[m]}\lambda_{i}^{k}\nabla f_{i}(x_{k}))^{\top}\left(x_{k}-x_{k+1}\right)$
$=d_{k}^{\top}B_{k}s_{k}$
$\leqslant\frac{b}{\alpha_{1}}\|s_{k}\|^{2}\ \ \ \ \forall k\geqslant 0,$ (4.24)
where $\lambda_{k}={{\left(\lambda_{1}^{k},\lambda_{2}^{k},\cdots,\lambda_{m}^{k}\right)}^{\top}}$is the solution of (3.9). In the convex case, we can observe that $\eta_{k}>0$ and by Theorem 4.3 get
$\|s_{k}\|=\|x_{k+1}-x_{k}\|\to 0,\ k\to\infty.$
Then,
@@
$m_{k}$ $=\max\{-\eta_{k},0\}+\sum\limits_{i\in[m]}\lambda_{i}^{k}(f_{i}(x_{k})-f_{i}(x_{k+1}))$
$=\sum\limits_{i\in[m]}\lambda_{i}^{k}(f_{i}(x_{k})-f_{i}(x_{k+1}))\to 0,\ k\to\infty.$
in (4.5). Based on the analysis results above, for a sufficiently large $k$, $B_{k+1}s_{k}\approx y_{k}$. Then (4.8) can be seen as an approximation of the secant equation in [3]. Remarkably, for some convex functions, both MFQNMO and MQNMO in [3] may exhibit comparable convergence rates. This observation can be supported by the numerical experiments presented in the following section. It can be seen that MFQNMO not only ensure the effectiveness of the algorithm for convex problems, but also guarantee the global convergence for nonconvex MOPs.
@@
$f_{\lambda}^{k}(x)=\sum\limits_{i\in[m]}\lambda_{i}^{k}f_{i}(x),$
@@
where $\lambda_{k}={{\left(\lambda_{1}^{k},\lambda_{2}^{k},\cdots,\lambda_{m}^{k}\right)}^{\top}}$is the solution of (3.9). Let us define
$\bar{s}_{k}=\nabla^{2}f_{\lambda}^{k}(x^{*})^{\frac{1}{2}}s_{k},\quad\bar{y}_{k}=\nabla^{2}f_{\lambda}^{k}(x^{*})^{-\frac{1}{2}}y_{k},\quad\bar{\gamma}_{k}=\nabla^{2}f_{\lambda}^{k}(x^{*})^{-\frac{1}{2}}\gamma_{k},$
$\bar{B}_{k}=\nabla^{2}f_{\lambda}^{k}(x^{*})^{-\frac{1}{2}}B_{k}\nabla^{2}f_{\lambda}^{k}(x^{*})^{-\frac{1}{2}},$ (4.26)
@@
Suppose that Assumption 4.1 and Assumption 4.3 hold. Let $\{x_{k}\}$ be the sequence generated by Algorithm 1. Then
$\lim\limits_{k\to\infty}\frac{\|(B_{k}-\nabla^{2}f_{\lambda}^{k}(x^{*}))s_{k}\|}{\|s_{k}\|}=0.$
@@
$(\bar{G}_{\lambda}^{k}-\nabla^{2}f_{\lambda}^{k}(x^{*}))s_{k}=y_{k}-\nabla^{2}f_{\lambda}^{k}(x^{*})s_{k},\quad k\geqslant 0.$ (4.28)
It follows by Assumption 4.3 (iii), (4.28) and the definition of $\bar{y}_{k}$, $\bar{s}_{k}$ that
$\|\bar{y}_{k}-\bar{s}_{k}\|$ $=\|\nabla^{2}f_{\lambda}^{k}(x^{*})^{-\frac{1}{2}}(\bar{G}_{\lambda}^{k}-\nabla^{2}f_{\lambda}^{k}(x^{*}))\nabla^{2}f_{\lambda}^{k}(x^{*})^{-\frac{1}{2}}\bar{s}_{k}\|$
$\leqslant\|\nabla^{2}f_{\lambda}^{k}(x^{*})^{-\frac{1}{2}}\|^{2}\|\bar{s}_{k}\|\|\bar{G}_{\lambda}^{k}-\nabla^{2}f_{\lambda}^{k}(x^{*})\|$
@@
$\leqslant\bar{c}\|\bar{s}_{k}\|\varepsilon_{k},$ (4.29)
where $\bar{c}=M\|\nabla^{2}f_{\lambda}^{k}(x^{*})^{-\frac{1}{2}}\|^{2}$, $\varepsilon_{k}=\max\{\|x_{k+1}-x^{*}\|,\|x_{k}-x^{*}\|\}.$ Since
$|\|\bar{y}_{k}\|-\|\bar{s}_{k}\||\leqslant\|\bar{y}_{k}-\bar{s}_{k}\|,$